OpenStudy (fuzzyball):

FAN AND MEDAL PLS HELP! (6 points) Breanna is planning to enter a bike race to support her favorite local charity. For each description, write an equation that could be used to solve the problem and then solve the problem. (a) Breanna’s sponsor plans to give $15 to her favorite charity for each mile she bikes. How many miles would Breanna need to bike if she wants to earn $450 for her favorite charity? Equation: Solution:

OpenStudy (fuzzyball):

(b) Breanna predicts that she may only be able to bike 25 mi. She still wants to earn $450 for her favorite charity. How much money should she ask her sponsor to give her per mile in order to reach her goal? Equation: Solution:

OpenStudy (noseboy908):

Is the equation then asking "How much should she ask for per mile to get $450 by only 25 miles?

OpenStudy (fuzzyball):

it is asking how many miles she need to ride to get $450 at $15 a mile

OpenStudy (noseboy908):

Ok, then let's find that first. Do you have any idea how we'd set up an equation to solve for that?

OpenStudy (fuzzyball):

450/15?

OpenStudy (noseboy908):

Yeah, solve that.

OpenStudy (fuzzyball):

as the Equation?

OpenStudy (fuzzyball):

wich would be 30 but how would i do the soulution?

OpenStudy (noseboy908):

Well, the important thing is to determine /what/ exactly the question is asking. Does it want to know how many miles she'd have to bike to make at least $450 at $15/mi?

OpenStudy (fuzzyball):

yea i think

OpenStudy (noseboy908):

My mistake, I misread it. You're correct. So let's create an equation for the problem.

OpenStudy (fuzzyball):

the equation would be 450/15=30? right???

OpenStudy (noseboy908):

That's how you would solve for the equation. Just work backwards a little. Try to incorporate a variable, x, into it.

OpenStudy (fuzzyball):

450/15=x?

OpenStudy (noseboy908):

Almost there! Now, work backwards just one more step! Isolate $450 please.

OpenStudy (fuzzyball):

15/450=X i dont know??? how??

OpenStudy (noseboy908):

Don't worry, just one step at a time. So you know that as an end result, you want the variable and the $15/mi to end up totaling $450, so we set that up on it's own, as (something) = $450. Now, let's add what we know. We're also told that they're willing to donate $15 PER mile, so what is the only unknown thing so far?

OpenStudy (fuzzyball):

how may miles

OpenStudy (noseboy908):

Yes! So let's set that as x. Thus we have our formula: (Price per mile) (quantity of miles) = (Total money made). Fill in what you can, please. Use x as the variable for what you don't know.

OpenStudy (fuzzyball):

15*X=450??

OpenStudy (noseboy908):

Perfect! Now, there's your first answer! That's your equation. Now, earlier you were eager to solve that equation for x, so go ahead and do that, please.

OpenStudy (fuzzyball):

450/15=30

OpenStudy (noseboy908):

Excellent! That's part A's equation and solution. Now, on to part B!

OpenStudy (noseboy908):

What, pray tell, does part B ask of us?

OpenStudy (fuzzyball):

25*X=450? she thinks she can only ride 25miles so she want to know how much her charity would have to pay her per mile if she rode 25 miles

OpenStudy (fuzzyball):

wich would be 18

OpenStudy (fuzzyball):

per mile they would have to pay her

OpenStudy (noseboy908):

Excellent work.

OpenStudy (noseboy908):

That is, then, the equation and solution for part B.

OpenStudy (fuzzyball):

thanks so much so last question

OpenStudy (noseboy908):

Fire away!

OpenStudy (fuzzyball):

so how would i do the soultution is numbers

OpenStudy (noseboy908):

The solution for A, assuming your teacher wants it in this form, would be x = 30 miles. For B, it'd be x = $18.00. Now, they may want you to elaborate too, so you'd have to verify how they want the solution presented to be sure.

OpenStudy (fuzzyball):

ok thanks so much you have been a HUGE help!

OpenStudy (noseboy908):

You are most welcome! Now, I only ask one final thing of you. By clicking on my name you can view my profile. From there, you can view testimonials. I'd greatly appreciate you leaving one for me. Be honest, and speak your mind.

OpenStudy (noseboy908):

Now, if you have any further questions, I'll be on later tonight to help. Just message me or mention me, and I'll help as soon as I see the notification. Au revoir!